OpenStudy (anonymous):

On what intervals within [0, 2π) is f increasing? f(x) = sin (x + π/6)

OpenStudy (anonymous):

find f'(x) > 0

OpenStudy (tkhunny):

Perhaps between the min and the max?

OpenStudy (anonymous):

I was thinking somewhere between [0, 5pi/12] U [11pi/12, 11pi/6] but its not working.. So i'm doing something wrong, but I don't know what.

OpenStudy (jdoe0001):

hmmm do you know the graph of sin(x) ?

OpenStudy (anonymous):

for f(x) to be increasing,f'(x)>0 \[f \prime \left( x \right)=\cos \left( x+\frac{ \pi }{ 6 } \right)>0\] \[\cos \left( x+\frac{ \pi }{ 6 } \right)\ge 0,~or~x+\frac{ \pi }{6 }\in \left( 0,\frac{ \pi }{ 2 } \right)\cup \left( \frac{ 3\pi }{2 },2\pi \ \right)\] \[x \in \left( \frac{ -\pi }{6 },\frac{ \pi }{2 }-\frac{ \pi }{ 6 } \right)\cup \left( \frac{ 3\pi }{ 2 }-\frac{ \pi }{6 },\left( 2\pi-\frac{ \pi }{6 } \right) \right)\] \[x \in \left( \frac{ -\pi }{ 6 },\frac{ \pi }{ 3 } \right)\cup \left(\frac{ 4\pi }{ 3 },\frac{ 11\pi }{ 6 } \right)\]

OpenStudy (jdoe0001):

\(\bf \textit{parent function}\implies f(x)=sin(x)\\ \quad \\ \begin{array}{llll} f(x)=sin&\left(x{\color{red}{ +\frac{\pi}{6}}}\right)\\ &\qquad \uparrow \\ &\textit{horizontal shift to the left} \end{array}\)

OpenStudy (anonymous):

Yea i know the graph of sin(x) but since this is like a phase shift problem doesnt that move everything over to the left. Because its a positive "c" value. like everything shifts by -pi/6 to the left. and since its between [0,2pi] would i still include the point [0, 5pi/12] as one of the increasing intervals? I get what it is supposed to do, but i somewhere im doing something wrong as far as actually figuring out what the intervals are supposed to be.

OpenStudy (anonymous):

I really appreciate everyone's feedback though. It has all been very helpful!

OpenStudy (mathmale):

May I assume you want to continue with the solution of this problem?

OpenStudy (anonymous):

Yes haha, the reply from surjithayer was very close, but it has to be between [0,2pi) and it won't work if it isnt. I tried using 0 instead of what he had, but it still wasnt working. I honestly don't know what i am doing wrong at this point

OpenStudy (mathmale):

Sirji has determined that the derivative of f(x)=sin(x + pi/6) is simply f '(x)=cos(x+pi/6); I agree. As he indicated, we need to determine on which intervals cos(x+pi/6) > 0, which is/are where f(x) is increasing. Use the sum formula for cos (a + b) = cos a cos b - sin a sin b to obtain cos (x+pi/6) = cos x cos pi/6 - sin x sin pi/6, or \[\cos (x+\frac{ \pi }{ 6 })=(\cos x)(\frac{ \sqrt{3} }{ 2 }) - (\sin x)(\frac{ 1 }{ 2 }) > 0\] Solving this inequality will lead to identification of the intervals on which the original function is increasing.

OpenStudy (anonymous):

when we find the interval as mr.mathmale suggested ,we can not divide by cos x or sin x ,because we are not clear they are positive or negative and negative sign changes the inequality. In my opinion \[x \in \left( 0,\frac{ \pi }{ 3 } \right)\cup \left( \frac{ 4\pi }{ 3 } ,\frac{ 11\pi }{ 6 }\right)\]